a stone is thrown upwards with an initial speed of 10m/s. calculate the time after which it acquires the same initial speed but travels in opposite direction.

Grade:10

2 Answers

vikas askiitian expert
13 years ago

total time of flight is the time after which it attains same speed but in

opposite direction...

T = 2(u/g)1/2 = 2sec

iit jee
13 years ago

let 'u' be initial velocity=10m/s.

we know that v=u-gt1(t1 is time of ascend, v=0,g=10m/s2)

t1=1 second.

during ascend u=0 and let t2 be time of descend,v=10m/s

by plugging the values in equation v=u+gt2

we get t2=1second

therefore the time after which it aquires same speed is t1+t2=2 seconds

plzzzzzzzz approve!!!!!!!!!!!!!!!